5(g-7)+2[g-3(g-5)]=0
Answer by jsmallt9(3758)

5(g-7)+2[g-3(g-5)]=0
Since subtractions are a source of many errors and since addition is commutative and associative, I recommend starting with rewriting all the subtractions as additions of the opposite:
5(g+(-7))+2[g+(-3)(g+(-5))]=0

Next we'll use the order of operations (aka PEMDAS) to simplify the left side. (The right side is already simplified.) Using the Distributive Property on (-3)(g+(-5)) we get:
5(g+(-7))+2[g+(-3g)+15]=0
Adding the like terms in the brackets, g and (-3g), we get:
5(g+(-7))+2[(-2g)+15]=0
Using the Distributive Property again (twice) we get:
5g + (-35) + (-4g) + 30 = 0
Adding like terms we get:
g + (-5) = 0

Now that the left side is simplified we can see that the equation is very simple to solve. Just add 5 to each side:
g = 5